Please Help:
Find the present value of the future amount. Assume 365 days in a year. Round to the nearest cent.
$23,749 for 107 days; money earns 8.5%
A) $577.39
B) $23,176.88
C) $21,888.48
D) $23,171.61

Answer by rfer(16322)

first find the interest rate for 107 days.
.085/365=.000232876*107=.024917808
23749/1.0249170=23171.61 (D)